=== 2019/6/6(Thur.) ===

==== Schedule ====
 * --(5/22-23: MZ release from PSL)--
 * --(5/20-24: Exchanging cryo valves.)--
 * --(5/27-6/2: Y arm mass lock.)-- (not done yet) 
 * --(5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS) )--
 * 6/3-6/14: Installation of TMS for ETMY
 * 6/8: Engineering run

 * ITMX
  * --(5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm)--
  * --(5/22AM~: tuned off shield cooler, finesse measurement for Xarm)--

 * ETMX
  * --(5/2: Cooling start)--

 * ITMY
  * --(5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm)--

 * ETMY
  * --(5/17~: tuned off ETMY cryostat cooler)--

 * Long term:
  * 4/22-5/31: Yarm commissioning	
  * 5/27-6/30: DRMI commissioning
  * 7/1-8/9: DRFPMI commissioning
  * 8/12-10/4: DRFPMI ASC
  * 10/7-11/15: Post Commissioning
  * 11/18-12/13: Noise hunting


==== Note ====
 * TMS installation this week.
 * Please reduce the number or speed of 16kHz channels.

==== Night work ====

==== Weekend ====

==== Report of work ====
 * COM: Finesse of X arm was OK with 1440.
 * COM: PRMI was locked. It was played with MICH sign flip to change resonant conditions between carrier and sidebands.
 * DGS: Maintenance was done 
 * AOS: Moved staff to proper position. Needs lights soon, this afternoon if the fastest case.
 * DGS: Installed BIO cards for LSC.
 * DGS: Dolphin issue happened at k1src.
 * DGS, AOS: A new rack was set at ETMY area. It will move in future to make more space for the road in front of racks.
 * PEM: Ground difference made a spark between 100V and 200V AC at 1500m area.
 * PEM: Reduced 16kHz channels.


==== Work plan ====